OpenStudy (anonymous):

HELP INTEGRATE :

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ x ^{2} }{ \sqrt[3]{x+1} }\]

OpenStudy (anonymous):

u = x+1 x = u-1 dx = du substitute, expand, power rule, back sub

OpenStudy (anonymous):

let me try that for a moment, i will be back if i fail

OpenStudy (anonymous):

THANK YOU @SOURWING. I HAVE GOT THE ANSWER NOW!